OpenStudy (anonymous):

Add. Simplify by collecting like radical terms if possible, assuming all expressions under radicals represent nonnegative numbers. sqrt3a +8sqrt12a^3

OpenStudy (anonymous):

\[\sqrt{3a} + 8\sqrt{12a ^{3}}\] \[\sqrt{3a}+ 8\sqrt{4*3*a*a*a}\] \[\sqrt{3a} + 16a \sqrt{3a}\] \[(1+16a)\sqrt{3a}\]